OpenStudy (driftracer305):

let u = (-3,4)........... write resulting vector in component form and find magnitude....

OpenStudy (driftracer305):

condition is 3u

OpenStudy (shamim):

-3 is x component of ur given vector u

OpenStudy (driftracer305):

yes.........i know....

OpenStudy (shamim):

4 is y component of vector u

OpenStudy (driftracer305):

i just want to know how to find magnitude.......... no clue about that

OpenStudy (shamim):

good

OpenStudy (zzr0ck3r):

\(||-3i+4j||=\sqrt{(-3)^2+4^2}\)

OpenStudy (shamim):

(-3)^2=?

OpenStudy (zzr0ck3r):

(-3)^2=3^2

OpenStudy (zzr0ck3r):

note: this is just the Pythagorean theorem.

OpenStudy (driftracer305):

k @zzr0ck3r ...... but equals 5......

OpenStudy (driftracer305):

so that's the magnitude...?

OpenStudy (zzr0ck3r):

\(||<-3,4>||=\sqrt{3^2+4^2}=\sqrt{25}=5\) Correct \(\huge \checkmark\)

OpenStudy (zzr0ck3r):

yes

OpenStudy (driftracer305):

lol......... dude... the answers says 15....... i am confused as well....

OpenStudy (zzr0ck3r):

were you given another vector?

OpenStudy (zzr0ck3r):

the fact that it says resultant vector makes me believe there was more than one vector and you were supposed to do something with it...

OpenStudy (zzr0ck3r):

surely the magnitude of <-3,4> is not 15

OpenStudy (driftracer305):

well it had v = (2,4)......... ohh m guessing it has something to do with this

OpenStudy (zzr0ck3r):

yes

OpenStudy (driftracer305):

lol....... sorry my bad

OpenStudy (zzr0ck3r):

post the whole question if you want the whole answer:P

OpenStudy (driftracer305):

k hang on a sec......

OpenStudy (driftracer305):

Let u = (-3,4), v = (2,4) and w = (4,-1)........Write each resulting vector in component form and find the magnitude............ question:- 3u @zzr0ck3r

OpenStudy (zzr0ck3r):

but either way, \(||<a_0,a_1,a_2,....,a_n>||=\sqrt{a_0^2+a_1^2+a_2^2+.....+a_n^2}\)

OpenStudy (driftracer305):

k si i knw i got vector correct as (-9, 12)

OpenStudy (zzr0ck3r):

still not writing the whole question.....

OpenStudy (zzr0ck3r):

resultant of what?

OpenStudy (driftracer305):

3u.....

OpenStudy (zzr0ck3r):

ok then magnitude is then \(\sqrt{9^2+12^2}=\sqrt{81+144}=\sqrt{225}=15\)

OpenStudy (zzr0ck3r):

and we win \(\huge \checkmark\)

OpenStudy (driftracer305):

ahh.............. dang it so simple .. so wat did v have to do with this thou??

OpenStudy (zzr0ck3r):

no idea

OpenStudy (zzr0ck3r):

are there more questions?

OpenStudy (driftracer305):

lol...... see that y i only put the value of u up there in first place.............. but yea there are more question..... but i think i am confident enough now to solve them on my own....................... thx to ur help.........lol.............. Thanks a ton man....... @zzr0ck3r

OpenStudy (zzr0ck3r):

npz